# ChunkKV: Semantic-Preserving KV Cache Compression for Efficient Long-Context LLM Inference

### v3

@@ Abstract
Large Language Models (LLMs) require significant GPU memory when processing long texts, with the key value (KV) cache consuming up to 70% of total memory during inference. Although existing compression methods reduce memory by evaluating the importance of individual tokens, they overlook critical semantic relationships between tokens, resulting in fragmented context and degraded performance. We introduce ChunkKV, which fundamentally reimagines KV cache compression by treating semantic chunks - rather than isolated tokens - as basic compression units. This approach preserves complete linguistic structures and contextual integrity, ensuring that essential meaning is retained even under aggressive compression. Our innovation includes a novel layer-wise index reuse technique that exploits the higher cross-layer similarity of preserved indices in ChunkKV, reducing computational overhead and improving throughput by 26.5%. Comprehensive evaluations on challenging benchmarks: LongBench, Needle-In-A-HayStack, GSM8K, and JailbreakV demonstrate that ChunkKV outperforms state-of-the-art methods by up to 8.7% in precision while maintaining the same compression ratio. These results confirm that semantic-aware compression significantly enhances both efficiency and performance for long-context LLM inference, providing a simple yet effective solution to the memory bottleneck problem.
**footnotetext: Equal Contribution.$\dagger$\$$\dagger$\$footnotetext: Corresponding Author.
@@ 1 Introduction
Large Language Models (LLMs) have become essential for addressing various downstream tasks of natural language processing (NLP), including summarization and question answering, which require the interpretation of a long context from sources such as books, reports, and documents, often encompassing tens of thousands of tokens [1, 2, 3]. Recent advances in long-context technology within the field of machine learning (ML) systems [4, 5, 6] have significantly improved computational throughputs and reduced latency of LLMs to process increasingly large input context lengths [7, 8] with historical KV cache (key value attentions). However, the memory requirement of the KV cache in serving super-long contexts becomes a new bottleneck [9, 10]. For instance, the KV cache for a single token in a 7B-parameter model requires approximately 0.5 MB of GPU memory, resulting in a 10,000-token prompt consuming around 5 GB of GPU memory.
@@
To address the substantial GPU memory consumption caused by KV caching, recent studies consider compressing the KV cache by pruning non-important discrete parts from the prompt tokens [9, 11, 12, 13, 14, 15, 16, 17]. H2O [9] and SnapKV [11] have shown that retaining less than 50% of the discrete KV cache can significantly reduce GPU memory usage with minimal impact on performance. However, we identify that the previous KV cache compression methods [9, 13] measure token importance isolatedly, neglecting the dependency between different tokens on the characteristics of the real-world language. For example, as shown in Figure 1, focusing on the importance of the token level can excessively focus on words about subjects “turaco” in the question while omitting crucial information about objects (foods) in the documents, resulting in the loss of essential semantic information. This motivates us to rethink the following question:
@@
In light of this, we observe that complete semantic information usually appears in a continuous sequence [18, 19, 20]. Thus, we introduce a straightforward yet effective ChunkKV, grouping the tokens in a chunk as a basic compressing unit, which should be preserved or discarded as a whole. Thus, it retains the most informative semantic chunks from the original KV cache. As shown in Figure 1, preserving a chunk helps catch the subject, predicate, and object. In Table 12, we quantify the minimal loss and higher recovery of the attention score achieved by the chunk-based approach at the inference stage. Furthermore, we investigate that the preserved KV cache indices by ChunkKV exhibit a higher similarity compared to previous methods. Consequently, we develop a technique called layer-wise index reuse, which reduces the additional computational time introduced by the KV cache compression method. As outlined in Table 1, recent highly relevant KV cache compression methods lack the ability to retain semantic information and efficiently reuse indices.
@@ 3.2 Chunk Based KV Compression
$\mathbf{A}\leftarrow\mathbf{Q}_{T_{q}-w:T_{q}}\mathbf{K}^{T}$triangleright Attention scores for the observe window
$C\leftarrow\left\lceil\frac{T_{k}}{c}\right\rceil$triangleright Calculate the number of chunks
@@
     $\mathbf{A}_{i}\leftarrow\sum_{j=(i-1)c+1}^{ic}\mathbf{A}_{:,j}$triangleright Sum of attention scores for each chunk
@@
[…]
@@
[…]
@@ 4 Experiment Results
In this section, we conduct experiments to evaluate the effectiveness of ChunkKV on KV cache compression in two benchmark fields, with a chunk size set to 10 even for various model architectures. The first is the In-Context Learning benchmark, for which we select GSM8K [23] and Jailbreakv [24] to evaluate the performance of ChunkKV, furthermore, we also include multi-step reasoning LLM DeepSeek-R1-Distill-Llama-8B [25] to evaluate the performance of ChunkKV. The In-Context Learning scenario is a crucial capability for LLMs and has been adapted in many powerful technologies such as Chain-of-Thought [30, 31, 32]. The second is the Long-Context benchmark, which includes LongBench [21] and Needle-In-A-HayStack (NIAH) [22], both widely used for assessing KV cache compression methods. All experiments were carried out three times, using the mean score to ensure robustness.
@@ 4.2 Long-Context Benchmark
KV cache Size SLM H2O SKV PKV ChunkKV (Ours)
@@ A.5 Implementation Details
3:$q_{\text{observe}}\leftarrow\mathbf{Q}_{T_{q}-w:T_{q}}$triangleright Extract observe window queries
4:$\mathbf{A}\leftarrow q_{\text{observe}}\mathbf{K}^{T}$triangleright $\mathbb{R}^{w\times T_{k}}$ attention scores
5:$C\leftarrow\left\lceil\frac{T_{k}}{c}\right\rceil$triangleright Number of chunks
@@
12:end for
13:$k\leftarrow\min(\left\lfloor\frac{L_{\text{max}}-w}{c}\right\rfloor,C)$triangleright Limit k to available chunks
@@
16:$\text{mask}\leftarrow\text{zeros}(T_{k})$triangleright Binary mask for token selection
@@
[…]
@@
25:$\text{indices}\leftarrow\text{nonzero}(\text{mask})$triangleright Get indices of tokens to keep
26:$\mathbf{K}^{\prime}\leftarrow\text{index\_select}(\mathbf{K},\text{indices})$
27:$\mathbf{V}^{\prime}\leftarrow\text{index\_select}(\mathbf{V},\text{indices})$
@@ B.4 Chunk Size
Table 18 and 19 show the performance of ChunkKV with different comperession ratios and different chunk sizes on the LongBench and NIAH. We conducted extensive experiments across different compression ratios and KV cache sizes to shows the effectiveness of ChunkKV and the chunk size is robust.
@@ B.6 KV Cache Quantization
B.7 Efficiency results
@@ Appendix C Theoretical Understanding
[…]
@@
[…]
@@
Thus, for $[S_{n},x_{\text{test}}]$, the in-context learning predictor can be written as $f_{n}(x_{\text{test}}):=\operatorname*{arg\,max}_{y}p(y|S_{n},x_{\text{test}})$, which outputs the most likely prediction over the pretraining distribution conditioned on the prompt distribution. Its expected 0-1 error with $n$ examples is $L_{\text{0-1}}(f_{n})=\mathbb{E}_{x_{\text{test}},y_{\text{test}}\sim p_{\text{prompt}}}[\mathbf{1}[f_{n}(x_{\text{test}})\neq y_{\text{test}}]]$.
@@
[…]
@@ Appendix D Additional Related Work
Layer-Wise Technique. The layer-wise technique is widely used in the training and inference of large language models (LLMs). LISA [49] is a layer-wise sampling method based on observations of the training dynamics of Low-Rank Adaptation (LoRA)[50] across layers. LAMB[51] is a layer-wise adaptive learning rate method that speeds up LLM training by stabilizing training convergence with large batch sizes. DoLa [52] employs layer-wise contrasting to reduce hallucinations during LLM inference.
@@
In parallel, retrieval benchmarks have largely relied on synthetic datasets, offering researchers precise control over variables such as the length of input tokens. This approach minimizes the impact of disparate parametric knowledge resulting from varied training methodologies. A significant body of recent work has concentrated on the development of synthetic tasks specifically for retrieval evaluation [22, 60, 61, 62, 63]. In addition, researchers have explored the potential of extended contexts in facilitating various forms of reasoning [64].
@@
KV Cache Quantization. Among quantization approaches for KV cache optimization, SmoothQuant [71] stands out as a notable post-training technique. By implementing balanced transformations between activation and weight quantization difficulties, this method enables 8-bit KV cache compression with minimal performance degradation. However, research by  [72] has demonstrated that SmoothQuant encounters substantial accuracy deterioration when compressed beyond 8 bits. In parallel developments, the FlexGen system [73] implements group-wise quantization at 4 bits for both key and value caches, while KIVI [37] introduces an innovative approach with asymmetric 2-bit per-channel quantization that requires no additional tuning phase.

### v4

@@ Abstract
Large Language Models (LLMs) require significant GPU memory when processing long texts, with the key value (KV) cache consuming up to 70% of total memory during inference. Although existing compression methods reduce memory by evaluating the importance of individual tokens, they overlook critical semantic relationships between tokens, resulting in fragmented context and degraded performance. We introduce ChunkKV, which fundamentally reimagines KV cache compression by treating semantic chunks - rather than isolated tokens - as basic compression units. This approach preserves complete linguistic structures and contextual integrity, ensuring that essential meaning is retained even under aggressive compression. Our innovation includes a novel layer-wise index reuse technique that exploits the higher cross-layer similarity of preserved indices in ChunkKV, reducing computational overhead and improving throughput by 26.5%. Comprehensive evaluations on challenging benchmarks: LongBench, Needle-In-A-HayStack, GSM8K, and JailbreakV demonstrate that ChunkKV outperforms state-of-the-art methods by up to 8.7% in precision while maintaining the same compression ratio. These results confirm that semantic-aware compression significantly enhances both efficiency and performance for long-context LLM inference, providing a simple yet effective solution to the memory bottleneck problem. The code is available at link.
[…]
@@ 1 Introduction
Large Language Models (LLMs) have become essential for addressing various downstream tasks of natural language processing (NLP), including summarization and question answering, which require the interpretation of a long context from sources such as books, reports, and documents, often encompassing tens of thousands of tokens [1, 2, 3, 4, 5]. Recent advances in long-context technology within the field of machine learning (ML) systems [6, 7, 8] have significantly improved computational throughputs and reduced latency of LLMs to process increasingly large input context lengths [9, 10] with historical KV cache (key value attentions). However, the memory requirement of the KV cache in serving super-long contexts becomes a new bottleneck [11, 12, 13, 14]. For instance, the KV cache for a single token in a 7B-parameter model requires approximately 0.5 MB of GPU memory, resulting in a 10,000-token prompt consuming around 5 GB of GPU memory.
@@
To address the substantial GPU memory consumption caused by KV caching, recent studies consider compressing the KV cache by pruning non-important discrete parts from the prompt tokens [11, 12, 15, 16, 17, 18, 19, 20, 21]. H2O [11] and SnapKV [15] have shown that retaining less than 50% of the discrete KV cache can significantly reduce GPU memory usage with minimal impact on performance. However, we identify that the previous KV cache compression methods [11, 17] measure token importance isolatedly, neglecting the dependency between different tokens on the characteristics of the real-world language. For example, as shown in Figure 1, focusing on the importance of the token level can excessively focus on words about subjects “turaco” in the question while omitting crucial information about objects (foods) in the documents, resulting in the loss of essential semantic information. This motivates us to rethink the following question:
@@
In light of this, we observe that complete semantic information usually appears in a continuous sequence [22, 4, 23, 24]. Thus, we introduce a straightforward yet effective ChunkKV, grouping the tokens in a chunk as a basic compressing unit, which should be preserved or discarded as a whole. Thus, it retains the most informative semantic chunks from the original KV cache. As shown in Figure 1, preserving a chunk helps catch the subject, predicate, and object. In Table 13, we quantify the minimal loss and higher recovery of the attention score achieved by the chunk-based approach at the inference stage. Furthermore, we investigate that the preserved KV cache indices by ChunkKV exhibit a higher similarity compared to previous methods. Consequently, we develop a technique called layer-wise index reuse, which reduces the additional computational time introduced by the KV cache compression method. As outlined in Table 1, recent highly relevant KV cache compression methods lack the ability to retain semantic information and efficiently reuse indices.
@@ 3.2 Chunk Based KV Compression
$\mathbf{A}\leftarrow\mathbf{Q}_{T_{q}-w:T_{q}}\mathbf{K}^{T}$ $\triangleright$ Attention scores for the observe window
$C\leftarrow\left\lceil\frac{T_{k}}{c}\right\rceil$ $\triangleright$ Calculate the number of chunks
@@
  $\mathbf{A}_{i}\leftarrow\sum_{j=(i-1)c+1}^{ic}\mathbf{A}_{:,j}$ $\triangleright$ Sum of attention scores for each chunk
@@
$\textit{Top\_K\_Indices}\leftarrow\text{indices of Top-}k\text{ chunks based on }\mathbf{A}_{i}$
@@
$\mathbf{K}^{\prime},\mathbf{V}^{\prime}\leftarrow\text{index\_select}(\mathbf{K},\mathbf{V},\textit{Top\_K\_Indices})$
@@ 4 Experiment Results
In this section, we conduct experiments to evaluate the effectiveness of ChunkKV on KV cache compression in two benchmark fields, with a chunk size set to 10 even for various model architectures. The first is the In-Context Learning benchmark, for which we select GSM8K [27] and Jailbreakv [28, 34, 35] to evaluate the performance of ChunkKV, furthermore, we also include multi-step reasoning LLM DeepSeek-R1-Distill-Llama-8B [29] to evaluate the performance of ChunkKV. The In-Context Learning scenario is a crucial capability for LLMs and has been adapted in many powerful technologies such as Chain-of-Thought [36, 37, 38, 39]. The second is the Long-Context benchmark, which includes LongBench [25] and Needle-In-A-HayStack (NIAH) [26], both widely used for assessing KV cache compression methods. All experiments were carried out three times, using the mean score to ensure robustness.
@@ 4.2 Long-Context Benchmark
KV cache Size
SLM
H2O SKV PKV
ChunkKV
@@ 4.5 Comparing with KV Quantization
4.6 Analysis of Hybrid Compression: Chunk-level vs. Token-level at Different Layer Depths
A critical question raised during the review process was whether the benefits of chunk-level compression diminish in deeper Transformer layers, where semantic information becomes more abstract and diffused. To investigate this, we designed a hybrid compression model that applies different strategies at varying network depths.
Experimental Setup
We created a hybrid version of the LLaMA-3-8B-Instruct model. For the first 16 layers (bottom half), we applied our chunk-based ChunkKV. For the final 16 layers (top half), we applied SnapKV, a state-of-the-art token-level compression method. We then compared this hybrid model’s performance on the diverse LongBench benchmark against pure ChunkKV and pure SnapKV at identical compression ratios.
Table 12: Performance of the hybrid compression model on LongBench. While the hybrid model shows strengths in global understanding tasks, pure ChunkKV achieves the best overall performance, validating the robustness of preserving semantic chunks even in deep layers.
Method Single-Doc QA Multi-Doc QA Summarization Few-shot Synthetic Code Avg. Score $\uparrow$
FullKV 32.19 34.59 24.96 68.48 36.96 54.41 41.46
SnapKV (10% Ratio) 28.11 32.55 24.12 67.81 36.01 55.67 40.15
Hybrid Model (10% Ratio) 28.38 30.37 24.54 67.87 36.37 55.32 39.80
ChunkKV (10% Ratio) 28.50 33.46 22.20 67.62 37.47 58.98 40.51
Analysis and Insights
The results in Table 12 provide two key insights. First, the pure ChunkKV model achieves the highest overall average score. This empirically validates our core hypothesis: preserving local semantic integrity via chunking is a robust and effective strategy across all layers of the model, even where information is highly processed.
Second, the hybrid model reveals a fascinating, task-dependent performance trade-off.
  • For local information retrieval tasks (e.g., Single- and Multi-Document QA), pure ChunkKV is significantly superior. These tasks often require retrieving precise, intact text fragments. ChunkKV’s ability to preserve complete linguistic units prevents critical information from being fragmented, which is essential in these scenarios.
  • For global understanding tasks (e.g., Summarization and Few-shot Learning), the hybrid model performs best. In these tasks, synthesizing information from across the entire context is key. In the deeper layers, where abstract representations are formed, the token-level SnapKV method may be more adept at retaining a broader, more diffuse set of globally important signals.
This analysis does not undermine our approach but rather enriches it, suggesting that the future of KV cache compression may lie in adaptive, task-aware strategies. However, for a general-purpose and robust solution, pure ChunkKV proves to be the most effective.
@@ 5 Conclusion
Acknowledge
This work was supported by the NSFC grant 62432008; RGC RIF grant R6021-20; an RGC TRS grant T43-513/23N-2; RGC CRF grants C7004-22G, C1029-22G and C6015-23G; NSFC/RGC grant CRS_HKUST601/24 and RGC GRF grants 16207922, 16207423 and 16203824; National Natural Science Foundation of China (Grant No.62506318); Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality;the Guangzhou Municipal Joint Funding Project with Universities and Enterprises under Grant No. 2024A03J0616 and Guangzhou Municipality Big Data Intelligence Key Lab (2023A03J0012).
Appendix
@@ A.5 Implementation Details
[…]
6:Allocate $\text{chunk\_scores}\in\mathbb{R}^{C}$ to store chunk importance
@@
11:  $\text{chunk\_scores}[i]\leftarrow\sum_{j=\text{start}}^{\text{end}-1}\sum_{q=0}^{w-1}\mathbf{A}_{q,j}$ $\triangleright$ Vectorized sum
12:end for
@@
[…]
@@
20:  $\text{mask}[\text{start}:\text{end}]\leftarrow 1$ $\triangleright$ Mark entire chunk as kept
@@
[…]
@@ B.4 Chunk Size
Table 19 and 20 show the performance of ChunkKV with different comperession ratios and different chunk sizes on the LongBench and NIAH. We conducted extensive experiments across different compression ratios and KV cache sizes to shows the effectiveness of ChunkKV and the chunk size is robust.
@@ B.6 KV Cache Quantization
B.7 Comparison with Orthogonal and Training-Based Methods
To better position ChunkKV in the broader landscape of inference optimization, we conducted additional experiments comparing our training-free eviction method against Palu [45], a recent training-based compression method. The experiments were conducted on the LongBench benchmark with the LLaMA-3-8B-Instruct model.
Table 24: Performance comparison of ChunkKV with Palu (training-based) on the LongBench benchmark (LLaMA-3-8B-Instruct). ChunkKV demonstrates a superior performance-efficiency trade-off compared to training-based methods on this diverse benchmark and remains competitive with orthogonal quantization techniques.
Method Single-Doc QA Multi-Doc QA Summarization Few-shot Synthetic Code Average Score $\uparrow$
FullKV (FP16) 32.19 34.59 24.96 68.48 36.96 54.41 41.46
Eviction Methods (Ours)
ChunkKV (10% Ratio) 28.50 33.46 22.20 67.62 37.47 58.98 40.51
ChunkKV (30% Ratio) 31.48 33.27 24.53 68.34 37.30 59.02 41.59
Training-Based Methods
Palu (50% Ratio) 8.77 7.43 20.71 61.62 8.59 18.43 22.48
Palu (70% Ratio) 10.06 8.31 23.71 68.64 35.00 38.97 28.84
The results in Table 24 highlight key differences. Palu, as a training-based method, experiences significant performance degradation on the diverse LongBench benchmark, which represents an Out-Of-Distribution (OOD) challenge relative to its training data. In contrast, our training-free method, ChunkKV, demonstrates robust performance.
B.8 Efficiency results
@@
Table 26: Latency and throughput comparison on LLaMA-3-8B-Instruct, including long-context scenarios up to 16k tokens. Percentages in parentheses indicate improvements over the FullKV baseline. ChunkKV with layer-wise reuse (ChunkKV_reuse) consistently delivers the best efficiency gains, especially in long-context settings.
Method Sequence Length Performance Metrics
Input Output Latency(s) $\downarrow$ Throughput(T/S) $\uparrow$
4k Context Length
FullKV 4096 1024 43.60 105.92
SnapKV 4096 1024 37.92 (13.0%) 120.42 (13.7%)
ChunkKV 4096 1024 37.52 (13.9%) 118.85 (12.2%)
ChunkKV_reuse 4096 1024 37.35 (14.3%) 124.09 (17.2%)
16k Context Length
FullKV 16384 1024 49.60 323.24
SnapKV 16384 1024 39.20 (21.0%) 381.11 (17.9%)
ChunkKV 16384 1024 38.82 (21.7%) 381.61 (18.1%)
ChunkKV_reuse 16384 1024 36.96 (25.5%) 389.21 (20.4%)
@@ Appendix C Theoretical Understanding
Pretraining Data Distribution. Given a set of concepts $\Theta$ and a concept $\theta\in\Theta$, we define the pretraining data is sampled from $p(o_{1},\dots,o_{T})=\int_{\theta\in\Theta}p(o_{1},\dots,o_{T}|\theta)p(\theta)d\theta$ [24]. Each token $o$ is sampled from a vocabulary $\mathbb{O}$}. For simplicity, we write $o_{1:t}=o_{1}\dots o_{t}$.
@@
Prompt Distribution. Following [24], a prompt is composed of an input token sequence $x$ followed by an output token $y$. Then, the $i$-th training example $\dagger$$\dagger\daggerHere$, training example in prompts means happens during the prompt learning. that can appear in any place in the whole prompt $o_{1:T}$ is defined as $O_{i}$ consisting of an input $x_{i}=O_{i}\left[1:k-1\right]$ (the first k1k-1 tokens) followed by the output $y_{i}=O_{i}\left[k\right]$ at the end, where the length $k$ is fixed for simplicity.
@@
Thus, for $S_{n},x_{\text{test}}]$, the in-context learning predictor can be written as $f_{n}(x_{\text{test}}):=\operatorname*{arg\,max}_{y}p(y|S_{n},x_{\text{test}})$, which outputs the most likely prediction over the pretraining distribution conditioned on the prompt distribution. Its expected 0-1 error with $n$ examples is $L_{\text{0-1}}(f_{n})=\mathbb{E}_{x_{\text{test}},y_{\text{test}}\sim p_{\text{prompt}}}[\mathbf{1}[f_{n}(x_{\text{test}})\neq y_{\text{test}}]]$.
@@
Token-level Importance Measure. The token-level KV cache methods usually calculate the importance of different tokens. Then, the KV cache with indexes that have higher importance will be preserved. Such indexes are normaly choosed as the attention score. Considering that the case in Figure 1, where each token in the $i$-th training $\dagger$$\dagger\daggerHere$, training means prompt learning [24]. example sequence ($O_{i}=[x_{i},y_{i}]$) might be compressed, and tokens are sparsified concretely without out dependency to other tokens. However, in the generation process of the $i$-th training example, $O_{i}=[x_{i},y_{i}]$ is sampled from $p(O_{i}|h^{\text{start}}_{i},\theta^{\star})$ and $p_{\theta}^{j}(o):=p(O[j]=o|O[1:j-1],\theta)$ of the $j$-th token with previous tokens and the analogous distribution $p^{j}_{prompt}:=p_{prompt}(O[j]=o|O[1:j-1])$. And the KL divergence is defined as $\text{KL}_{j}(\theta^{\star}||\theta):=\mathbb{E}_{O[1:j-1]\sim p_{prompt}}[\text{KL}(p^{j}_{prompt}||p_{\theta}^{j})]$, which means that in a training example $O_{i}=[x_{i},y_{i}]=O_{i}[1:k]$, each token $O_{i}[j]$ has strong dependency with $O_{i}[1:j-1]$, noises on previous any $j$-th token will influence the distinguishability on the following tokens, i.e. requiring larger $\left\{\text{KL}_{u}(\theta^{\star}||\theta)\right\}_{u>j}$.
@@ Appendix D Additional Related Work
Layer-Wise Technique. The layer-wise technique is widely used in the training and inference of large language models (LLMs). LISA [57] is a layer-wise sampling method based on observations of the training dynamics of Low-Rank Adaptation (LoRA)[58, 34, 35, 37, 59] across layers. LAMB[60] is a layer-wise adaptive learning rate method that speeds up LLM training by stabilizing training convergence with large batch sizes. DoLa [61] employs layer-wise contrasting to reduce hallucinations during LLM inference.
@@
In parallel, retrieval benchmarks have largely relied on synthetic datasets, offering researchers precise control over variables such as the length of input tokens. This approach minimizes the impact of disparate parametric knowledge resulting from varied training methodologies. A significant body of recent work has concentrated on the development of synthetic tasks specifically for retrieval evaluation [26, 69, 70, 71, 72]. In addition, researchers have explored the potential of extended contexts in facilitating various forms of reasoning [73]. Liu et al. [74] explores the influence of the KV cache method beyond long-context benchmarks. LongGenBench [75] advances beyond traditional benchmarks by redesigning the format of questions and necessitating that LLMs respond with a single, cohesive long-context answer.
@@
KV Cache Quantization. Among quantization approaches for KV cache optimization, SmoothQuant [82] stands out as a notable post-training technique. By implementing balanced transformations between activation and weight quantization difficulties, this method enables 8-bit KV cache compression with minimal performance degradation. However, research by  [83] has demonstrated that SmoothQuant encounters substantial accuracy deterioration when compressed beyond 8 bits. In parallel developments, the FlexGen system [84] implements group-wise quantization at 4 bits for both key and value caches, while KIVI [44] introduces an innovative approach with asymmetric 2-bit per-channel quantization that requires no additional tuning phase. Additionally, AnTKV [85] introduces a two-stage token-aware vector quantization framework, leveraging gradient-weighted centroid learning and anchor-token selection to achieve ultra-low-bit KV cache quantization with minimal accuracy loss.